# Distributed Optimal Coverage Control in Multi-agent Systems: Known and Unknown Environments

### v2

@@ II Notations and Preliminaries
[…]
@@
$\mathbf{L}_{i,j}=\begin{cases}-l_{i,j},&\text{for}\quad i\neq j,\\ \sum_{j=1}^{n}l_{i,j},&\text{for}\quad i=j.\\ \end{cases}$ (1)
@@ IV Proposed Method
Based on the definition of $R_{i}\big{(}\mathbf{q},\mathbf{P}(t)\big{)}$, each agent $i$ at time $t\geq 0$ assigns a rank $R_{i}\big{(}\mathbf{q},\mathbf{P}(t)\big{)}$ between 00 to $n-1$ to all the points in the environment, where the points that the closest agent is the $i$-th agent receive $(0)$-rank, i.e. $R_{i}\big{(}\mathbf{q},\mathbf{P}(t)\big{)}=0$, the points that the second closest agent is the $i$-th agent receive $(1)$-rank, i.e. $R_{i}\big{(}\mathbf{q},\mathbf{P}(t)\big{)}=1$, and the points that the farthest agent is the $i$-th agent receive at most $(n-1)$-rank, i.e. $R_{i}\big{(}\mathbf{q},\mathbf{P}(t)\big{)}=n-1$. It should be noted that points with the same distance from two agents, i.e., $\lVert\mathbf{q}-\mathbf{p}_{j}(t)\rVert=\lVert\mathbf{q}-\mathbf{p}_{i}(t)\rVert$, are assigned the same rank by each of those agents, which translates to $R_{i}=R_{j}$ for that particular point. These points lie on the perpendicular bisector of $\mathbf{p}_{i}$ and $\mathbf{p}_{j}$ in $\mathbf{Q}$. In this work, the following cost function is introduced:
@@
and $\lambda(t):\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ where $\lim_{t\to\infty}\lambda(t)=0$.
@@
$\frac{1}{2}{\sum_{j=1}^{n}}\int_{\mathbf{Q}}\frac{\partial h_{\lambda}\!(t,R_{j})}{\partial\mathbf{p}_{i}(t)}\lVert\mathbf{q}-\mathbf{p}_{i}(t)\rVert^{2}\phi(\mathbf{q})\mathrm{d}\mathbf{q}=0.$ (13)
Proof. Using (8) and chain rule, (13) can be written as
$\frac{1}{2}$ $\sum_{j=1}^{n}\int_{\mathbf{Q}}\bigg{[}\frac{\partial h_{\lambda}\!(t,R_{j})}{\partial R_{j}}\lVert\mathbf{q}-\mathbf{p}_{i}(t)\rVert^{2}\phi(\mathbf{q})$
@@
$-\frac{1}{2}$ $\sum_{j=1}^{n}\int_{\mathbf{Q}}\bigg{[}\frac{\partial h_{\lambda}\!(t,R_{j})}{\partial R_{j}}\lVert\mathbf{q}-\mathbf{p}_{i}(t)\rVert^{2}\phi(\mathbf{q})$ (14)
@@
$+$ $\sum_{j=1}^{n}\int_{\mathbf{Q}}\bigg{[}\frac{\partial h_{\lambda}\!(t,R_{j})}{\partial R_{j}}\lVert\mathbf{q}-\mathbf{p}_{i}(t)\rVert^{2}\phi(\mathbf{q})$ (15)
@@
The choice of the parameter $\lambda(t)$ can be of great importance. Exponential functions, such as $\lambda(t)=\lambda_{\mathrm{s}}\alpha^{-\lambda_{\mathrm{f}}t}$ where $\lambda_{\mathrm{s}},\lambda_{\mathrm{f}}>0$, and $\alpha>1$, are often considered due to their mathematical simplicity and their ability to model various types of dynamics.
@@
Unlike $h_{i}(\cdot)$ given by (10) which has a binary value, $h_{\lambda}(\cdot)$ can have any value in the range of $(0,1]$. According to Definition 8, the closest agents to the point $\mathbf{q}\in\mathbf{Q}$ gets the maximum value of $h_{\lambda}(\cdot)$, while the remaining agents receive $10<h_{\lambda}<$ 1.
@@
Considering (9) and (11), for any fixed configuration of agents $\mathbf{P}\triangleq\{\mathbf{p}_{1},\dots,\mathbf{p}_{n}\}$, the following equation holds:
@@ IV-A Known Environment
In this section, a known environment is considered and a control signal for each agent is designed such that the proposed cost function $\hat{H}(\cdot)$ introduced by (11) is minimized. For this purpose, similar to the Voronoi-based algorithms, the gradient descent method is used to update the position of each agent as follows:
@@
$\frac{\partial\hat{H}(t)}{\partial\mathbf{p}_{i}(t)}=$ $\frac{1}{2}\sum_{j=1}^{n}\int_{\mathbf{Q}}\bigg{[}h_{\lambda}\!(t,R_{i})\frac{\partial(\lVert\mathbf{q}-\mathbf{p}_{j}(t)\rVert^{2})}{\partial\mathbf{p}_{i}(t)}$
$\qquad+\frac{\partial h_{\lambda}\!(t,R_{i})}{\partial\mathbf{p}_{i}(t)}\lVert\mathbf{q}-\mathbf{p}_{j}(t)\rVert^{2}\Bigg{]}\phi(\mathbf{q})\mathrm{d}\mathbf{q}.$ (21)
@@
$\quad\qquad+\frac{1}{2}{\sum_{j=1}^{n}}\int_{\mathbf{Q}}\frac{\partial h_{\lambda}\!(t,R_{j})}{\partial\mathbf{p}_{i}(t)}\lVert\mathbf{q}-\mathbf{p}_{i}(t)\rVert^{2}\phi(\mathbf{q})\mathrm{d}\mathbf{q}.$ (22)
@@
$\mathbf{u}_{i}(t)=\epsilon W_{i}(t),$ (23)
@@
$\mathbf{W}_{i}(t)\triangleq\int_{\mathbf{Q}}h_{\lambda}\!(t,R_{i})\big{(}\mathbf{q}-\mathbf{p}_{i}(t)\big{)}\phi(\mathbf{q})\mathrm{d}\mathbf{q}.$ (24)
The following lemma gives some facts about $\mathbf{W}_{i}(t)$.
@@
$\mathbf{W}_{i}(t)$ and $\frac{\partial\mathbf{W}_{i}(t)}{\partial\mathbf{p}_{i}(t)}$ are bounded.
Proof. Since $\mathbf{Q}$ is a bounded environment, and $h_{\lambda}(\cdot)$ and $\phi(\cdot)$ are bounded, hence $\mathbf{W}_{i}(t)$ is also bounded. In addition, using the chain rule, we have
$\frac{\partial\mathbf{W}_{i}(t)}{\partial\mathbf{p}_{i}(t)}=$ $\int_{\mathbf{Q}}\Big{[}\frac{\partial h_{\lambda}\!(t,R_{i})}{\partial R_{i}}\ \!\frac{\partial R_{i}(\mathbf{q},\mathbf{P}(t))}{\partial\mathbf{p}_{i}(t)}\ \!(\mathbf{q}-\mathbf{p}_{i}(t))^{\mathrm{T}}$
$\qquad-h_{\lambda}\!(t,R_{i})\mathbf{I}_{N}\Big{]}\phi(\mathbf{q})\mathrm{d}\mathbf{q}.$ (25)
Since $h_{\lambda}\!(t,R_{i})\in(0,1]$, then the second term of (IV-A) is bounded as follows
$0<\int_{\mathbf{Q}}h_{\lambda}\!(t,R_{i})\phi(\mathbf{q})\mathrm{d}\mathbf{q}\leq\int_{\mathbf{Q}}\phi(\mathbf{q})\mathrm{d}\mathbf{q}<\eta.$ (26)
Now, regarding the first term of (IV-A), $\frac{\partial h_{\lambda}\!(t,R_{i})}{\partial R_{i}}$ is bounded according to the definition of $h_{\lambda}$. Moreover, $\frac{\partial R_{i}(\mathbf{q},\mathbf{P}(t))}{\partial\mathbf{p}_{i}(t)}$ is always zero except when $\lVert\mathbf{q}-\mathbf{p}_{i}(t)\rVert=\lVert\mathbf{q}-\mathbf{p}_{l}(t)\rVert$ since
$\frac{\partial R_{i}\big{(}\mathbf{q},\mathbf{P}(t)\big{)}}{\partial\mathbf{p}_{i}(t)}\!=\!$ ${\sum_{l=1}^{n}\frac{\partial\big{(}\lVert\mathbf{q}-\mathbf{p}_{i}(t)\rVert^{2}-\lVert\mathbf{q}-\mathbf{p}_{l}(t)\rVert^{2}\big{)}}{\partial\mathbf{p}_{i}(t)}}$
@@
$-2\sum_{l=1}^{n}\int_{\mathbf{Q}}\frac{\partial h_{\lambda}\!(t,R_{i})}{\partial R_{i}}\!\big{(}\mathbf{q}-\mathbf{p}_{i}(t)\big{)}\!\big{(}\mathbf{q}-\mathbf{p}_{i}(t)\big{)}^{\mathrm{T}}\phi(\mathbf{q})$
$\qquad\qquad\qquad\delta\big{(}\lVert\mathbf{q}-\mathbf{p}_{i}(t)\rVert^{2}-\lVert\mathbf{q}-\mathbf{p}_{l}(t)\rVert^{2}\big{)}\mathrm{d}\mathbf{q}$
$=-2\sum_{l=1}^{n}\int_{\mathcal{S}_{il}}\frac{\partial h_{\lambda}\!(t,R_{i})}{\partial R_{i}}\!\big{(}\mathbf{s}-\mathbf{p}_{i}(t)\big{)}\!\big{(}\mathbf{s}-\mathbf{p}_{i}(t)\big{)}^{\mathrm{T}}\phi(\mathbf{s})\mathrm{d}\mathbf{s},$ (28)
where (28) is the line integral over $\mathcal{S}_{il}\subset\mathbf{Q}$. Therefore, $\frac{\partial\mathbf{W}_{i}(t)}{\partial\mathbf{p}_{i}(t)}$ is bounded. $\blacksquare$
@@
$\dot{V}\big{(}t,\mathbf{P}(t)\big{)}$ $=\sum_{i=1}^{n}\frac{\partial\hat{H}^{\mathrm{T}}\big{(}t,\mathbf{P}(t)\big{)}}{\partial\mathbf{p}_{i}(t)}\ \!\dot{\mathbf{p}}_{i}(t),$ (30)
Using Lemma 3, (6) and (23), we can write
$\dot{V}\big{(}t,\mathbf{P}\big{)}=$ $-\epsilon\sum_{i=1}^{n}\mathbf{W}_{i}^{\mathrm{T}}(t)\mathbf{W}_{i}(t).$ (31)
Similarly, taking the second-order time derivative of $V(\cdot)$ yields:
$\ddot{V}\big{(}t,\mathbf{P}\big{)}$ $=-2\epsilon^{2}\sum_{i=1}^{n}\mathbf{W}_{i}^{\mathrm{T}}(t)\frac{\partial\mathbf{W}_{i}(t)}{\partial\mathbf{p}_{i}(t)}\mathbf{W}_{i}(t).$ (32)
According to Lemma 4, $\ddot{V}(\cdot)$ is bounded, and consequently $\dot{V}(\cdot)$ is uniformly continuous [10]. Since $V(\cdot)\geq 0$ (is lower bounded) and $\dot{V}(\cdot)\leq 0$, then $V(\cdot)$ has a finite limit. Therefore, using Barbalat’s lemma, $\dot{V}(\cdot)\to 0$. Considering (31) and (24), it follows that $\lim_{t\to\infty}u_{i}(t)=\lim_{t\to\infty}W_{i}(t)=0$ and agents converges to their optimal coverage configuration. $\blacksquare$

### v3

@@ II Notations and Preliminaries
In order to show the problem addressed in this paper, the conventional representation of required variables such as matrices and vectors is presented in bold. $\mathbb{R}$ and $\mathbb{R}_{\geq 0}$, respectively, stand for the real number field and the set of non-negative real numbers. The set of $n\times 1$ real column vectors is denoted by $\mathbb{R}^{n}$ and the set of $n\times m$ real matrices is denoted by $\mathbb{R}^{n\times m}$. Furthermore, $\mathbf{I}_{n}$ represents $n\times n$ identity matrix, $\mathbf{0}_{n\times m}$ represents $n\times m$ zero matrix, $\mathbf{1}_{n\times m}$ represents $n\times m$ one matrix, $(\cdot)^{\mathrm{T}}$ represents transpose, $|\cdot|$ is the absolute value, and $\lVert\cdot\rVert$ is the Euclidean norm.
@@
[…]
@@ IV Proposed Method
Based on the definition of $R_{i}\big{(}\mathbf{q},\mathbf{P}(t)\big{)}$, each agent $i$ at time $t\geq 0$ assigns a rank $R_{i}\big{(}\mathbf{q},\mathbf{P}(t)\big{)}$ between 0 to n1n-1 to all the points in the environment, where the points that the closest agent is the $i$-th agent receive $(0)$-rank, i.e. $R_{i}\big{(}\mathbf{q},\mathbf{P}(t)\big{)}=0$, the points that the second closest agent is the $i$-th agent receive $(1)$-rank, i.e. $R_{i}\big{(}\mathbf{q},\mathbf{P}(t)\big{)}=1$, and the points that the farthest agent is the $i$-th agent receive at most (n1)(n-1)-rank, i.e. $R_{i}\big{(}\mathbf{q},\mathbf{P}(t)\big{)}=n-1$. It should be noted that points with the same distance from two agents, i.e., $\lVert\mathbf{q}-\mathbf{p}_{j}(t)\rVert=\lVert\mathbf{q}-\mathbf{p}_{i}(t)\rVert$, are assigned the same rank by each of those agents, which translates to $R_{i}=R_{j}$ for that particular point. These points lie on the perpendicular bisector of $\mathbf{p}_{i}$ and $\mathbf{p}_{j}$ in $\mathbf{Q}$. In this work, the following cost function is introduced:
@@
and $\lambda(t):\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ where $\lim_{t\to\infty}\lambda(t)=0$. Note that the function $\lambda:[0,\infty)\to\mathbb{R}$ is continuously differentiable, strictly positive, strictly decreasing to zero ($\dot{\lambda}(t)<0,t\geq 0$) with a bounded second derivative, and satisfies $\lambda(0)=\lambda_{0}>0$, where $\lambda_{0}$ is finite.
@@
$\frac{1}{2}{\sum_{j=1}^{n}}\int_{\mathbf{Q}}\frac{\partial h_{\lambda}\!(t,R_{j})}{\partial\mathbf{p}_{i}(t)}\lVert\mathbf{q}-\mathbf{p}_{j}(t)\rVert^{2}\phi(\mathbf{q})\mathrm{d}\mathbf{q}=0.$ (13)
Proof. Using (8) and chain rule, left-side of (13) can be written as
$\frac{1}{2}$ $\sum_{j=1}^{n}\int_{\mathbf{Q}}\bigg{[}\frac{\partial h_{\lambda}\!(t,R_{j})}{\partial R_{j}}\lVert\mathbf{q}-\mathbf{p}_{j}(t)\rVert^{2}\phi(\mathbf{q})$
@@
$-\frac{1}{2}$ $\sum_{j=1}^{n}\int_{\mathbf{Q}}\bigg{[}\frac{\partial h_{\lambda}\!(t,R_{j})}{\partial R_{j}}\lVert\mathbf{q}-\mathbf{p}_{j}(t)\rVert^{2}\phi(\mathbf{q})$ (14)
@@
$+$ $\sum_{j=1}^{n}\int_{\mathbf{Q}}\bigg{[}\frac{\partial h_{\lambda}\!(t,R_{j})}{\partial R_{j}}\lVert\mathbf{q}-\mathbf{p}_{j}(t)\rVert^{2}\phi(\mathbf{q})$ (15)
@@
The choice of the parameter $\lambda(t)$ is arbitrary. However, exponential functions, such as $\lambda(t)=\lambda_{\mathrm{s}}\alpha^{-\lambda_{\mathrm{f}}t}$ where $\lambda_{\mathrm{s}},\lambda_{\mathrm{f}}>0$, and $\alpha>1$, are often considered due to their mathematical simplicity and their ability to model various types of dynamics.
@@
Unlike $h_{i}(\cdot)$ given by (10) which has a binary value, $h_{\lambda}(\cdot)$ can have any value in the range of $(0,1]$. According to Definition 8, the closest agents to the point $\mathbf{q}\in\mathbf{Q}$ gets the maximum value of $h_{\lambda}(\cdot)$, while the remaining agents receive $0<h_{\lambda}<1$.
@@
Considering (9) and (11), for any configuration of agents $\mathbf{P}\triangleq\{\mathbf{p}_{1},\dots,\mathbf{p}_{n}\}$, the following equation holds:
@@ IV-A Known Environment
[…]
@@
$\frac{\partial\hat{H}(t)}{\partial\mathbf{p}_{i}(t)}=$ $\frac{1}{2}\sum_{j=1}^{n}\int_{\mathbf{Q}}\bigg{[}h_{\lambda}\!(t,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}R_{j}})\frac{\partial(\lVert\mathbf{q}-\mathbf{p}_{j}(t)\rVert^{2})}{\partial\mathbf{p}_{i}(t)}$
$\qquad+\frac{\partial h_{\lambda}\!(t,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}R_{j}})}{\partial\mathbf{p}_{i}(t)}\lVert\mathbf{q}-\mathbf{p}_{j}(t)\rVert^{2}\Bigg{]}\phi(\mathbf{q})\mathrm{d}\mathbf{q}.$ (21)
@@
$\quad\qquad\quad+\frac{1}{2}{\sum_{j=1}^{n}}\int_{\mathbf{Q}}\frac{\partial h_{\lambda}\!(t,R_{j})}{\partial\mathbf{p}_{i}(t)}\lVert\mathbf{q}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{p}_{j}(t)}\rVert^{2}\phi(\mathbf{q})\mathrm{d}\mathbf{q}.$ (22)
@@
$\mathbf{u}_{i}(t)=\epsilon\mathbf{W}_{i}(t,\mathbf{P}(t)),$ (23)
@@
$\mathbf{W}_{i}(t,\mathbf{P}(t))\triangleq\int_{\mathbf{Q}}h_{\lambda}\!(t,R_{i})\big{(}\mathbf{q}-\mathbf{p}_{i}(t)\big{)}\phi(\mathbf{q})\mathrm{d}\mathbf{q}.$ (24)
The following lemma gives some facts about $\mathbf{W}_{i}(t,\mathbf{P}(t))$.
@@
$\mathbf{W}_{i}(t,\mathbf{P}(t))$ defined in (24) and $\frac{\partial\mathbf{W}_{i}(t)}{\partial\mathbf{p}_{j}(t)}$ are bounded for all $i,j\in\{1,\ldots,n\}$ and $t\geq 0$.
Proof. Since $\mathbf{Q}$ is a bounded environment, and, $h_{\lambda}(\cdot)$ and $\phi(\cdot)$ are bounded, $\mathbf{W}_{i}(t,\mathbf{P}(t))$ is also bounded. In addition, using the chain rule, we have
$\frac{\partial\mathbf{W}_{i}(t,\mathbf{P}(t))}{\partial\mathbf{p}_{j}(t)}=$ $\int_{\mathbf{Q}}\Big{[}\frac{\partial h_{\lambda}\!(t,R_{i})}{\partial R_{i}}\ \!\frac{\partial R_{i}(\mathbf{q},\mathbf{P}(t))}{\partial\mathbf{p}_{j}(t)}\ \!(\mathbf{q}-\mathbf{p}_{i}(t))^{\mathrm{T}}$
$\qquad-h_{\lambda}\!(t,R_{i})\frac{\partial\mathbf{p}_{i}}{\partial\mathbf{p}_{j}}\Big{]}\phi(\mathbf{q})\mathrm{d}\mathbf{q}.$ (25)
$\frac{\partial h_{\lambda}\!(t,R_{i})}{\partial R_{i}}$ is bounded according to the definition of $h_{\lambda}$. Moreover, $\frac{\partial R_{i}(\mathbf{q},\mathbf{P}(t))}{\partial\mathbf{p}_{j}(t)}$ is obtained as follows:
$\frac{\partial R_{i}\big{(}\mathbf{q},\mathbf{P}(t)\big{)}}{\partial\mathbf{p}_{j}(t)}\!=\!$ ${\sum_{l=1}^{n}\frac{\partial\big{(}\lVert\mathbf{q}-\mathbf{p}_{i}(t)\rVert^{2}-\lVert\mathbf{q}-\mathbf{p}_{l}(t)\rVert^{2}\big{)}}{\partial\mathbf{p}_{j}(t)}}$
@@
$\int_{\mathbf{Q}}f(\mathbf{q})\delta(g(\mathbf{q}))\mathrm{d}\mathbf{q}=\int_{\{\mathbf{q}\in\mathbf{Q}|g(\mathbf{q})=0\}}\frac{f(\mathbf{s})}{\|\nabla g(\mathbf{s})\|}\mathrm{d}\mathbf{s},$ (27)
we have for $j=i$:
$\frac{\partial\mathbf{W}_{i}}{\partial\mathbf{p}_{i}}=$ $\sum_{k\neq i}\int_{\mathcal{S}_{ik}}\frac{\partial h_{\lambda}\!(t,R_{i})}{\partial R_{i}}\frac{(\mathbf{s}-\mathbf{p}_{i})(\mathbf{s}-\mathbf{p}_{i})^{\mathrm{T}}\phi(\mathbf{s})}{\|\mathbf{p}_{k}-\mathbf{p}_{i}\|}\mathrm{d}\mathbf{s}$
$-\int_{\mathbf{Q}}h_{\lambda}(t,R_{i})\phi\mathrm{d}\mathbf{q}\mathbf{I}_{N},$ (28)
and for $j\neq i$:
$\frac{\partial\mathbf{W}_{i}}{\partial\mathbf{p}_{j}}=-\int_{\mathcal{S}_{ij}}\frac{\partial h_{\lambda}\!(t,R_{i})}{\partial R_{i}}\frac{(\mathbf{s}-\mathbf{p}_{j})(\mathbf{s}-\mathbf{p}_{i})^{\mathrm{T}}\phi(\mathbf{s})}{\|\mathbf{p}_{j}-\mathbf{p}_{i}\|}\mathrm{d}\mathbf{s}.$ (29)
Since $h_{\lambda}\!(t,R_{i})\in(0,1]$, then the second term of (28) is bounded as follows
$0<\int_{\mathbf{Q}}h_{\lambda}\!(t,R_{i})\phi(\mathbf{q})\mathrm{d}\mathbf{q}\leq\int_{\mathbf{Q}}\phi(\mathbf{q})\mathrm{d}\mathbf{q}<\eta,$ (30)
and this completes the proof. $\blacksquare$
@@
$\dot{V}\big{(}t,\mathbf{P}(t)\big{)}$ $=\frac{\partial\hat{H}}{\partial t}+\sum_{i=1}^{n}\frac{\partial\hat{H}^{\mathrm{T}}\big{(}t,\mathbf{P}(t)\big{)}}{\partial\mathbf{p}_{i}(t)}\ \!\dot{\mathbf{p}}_{i}(t).$ (32)
Using Lemma 3 and (23), we can write:
$\dot{V}\big{(}t,\mathbf{P}(t)\big{)}=$ $\frac{\partial\hat{H}}{\partial t}-\epsilon\sum_{i=1}^{n}\mathbf{W}_{i}^{\mathrm{T}}(t,\mathbf{P}(t))\mathbf{W}_{i}(t,\mathbf{P}(t)).$ (33)
The derivative $\frac{\partial\hat{H}}{\partial t}$ is computed as follows:
$\frac{\partial\hat{H}}{\partial t}$ $=\!-\frac{1}{2}\!\sum_{i=1}^{n}\!\int_{Q}\!\frac{\partial}{\partial t}\Big{(}\frac{R_{i}}{\lambda(t)}\Big{)}h_{\lambda}(t,R_{i})\|\mathbf{q}\!-\!\mathbf{p}_{i}(t)\|^{2}\!\phi\!\left(\mathbf{q}\right)\!\mathrm{d}\mathbf{q}$
$=\!\frac{1}{2}\!\sum_{i=1}^{n}\!\int_{Q}\!\frac{R_{i}\dot{\lambda}(t)}{\lambda^{2}(t)}h_{\lambda}(t,R_{i})\|\mathbf{q}\!-\!\mathbf{p}_{i}(t)\|^{2}\phi\!\left(\mathbf{q}\right)\!\mathrm{d}\mathbf{q}.$ (34)
Hence, (33) can be written as follows:
$\!\dot{V}\big{(}t,\mathbf{P}(t)\big{)}\!\!=$ $\frac{1}{2}\!\sum_{i=1}^{n}\!\int_{Q}\!\frac{R_{i}\dot{\lambda}(t)}{\lambda^{2}(t)}h_{\lambda}(t,R_{i})\|\mathbf{q}\!-\!\mathbf{p}_{i}(t)\|^{2}\phi\!\left(\mathbf{q}\right)\!\mathrm{d}\mathbf{q}$
$-\frac{1}{2}\sum_{i=1}^{n}2\epsilon\mathbf{W}^{\mathrm{T}}_{i}(t,\mathbf{P}(t))\mathbf{W}_{i}(t,\mathbf{P}(t)).$ (35)
Moreover, $\ddot{V}\big{(}t,\mathbf{P}(t)\big{)}$ is obtained as follows:
$\ddot{V}\big{(}t$ $,\mathbf{P}(t)\big{)}\!=$
$-\sum_{i=1}^{n}\!\Big{(}\!\int_{Q}\!\frac{R_{i}\dot{\lambda}(t)}{\lambda^{2}(t)}h_{\lambda}(t,R_{i})(\mathbf{q}\!-\!\mathbf{p}_{i}(t))\phi(\mathbf{q})\,\mathrm{d}\mathbf{q}\Big{)}^{\mathrm{T}}\dot{\mathbf{p}}_{i}(t)$
$\!+\!\frac{1}{2}\!\sum_{i=1}^{n}\!\int_{Q}\!\Big{(}\frac{R_{i}\dot{\lambda}(t)}{\lambda^{2}(t)}\Big{)}^{2}h_{\lambda}(t,R_{i})\|\mathbf{q}\!-\!\mathbf{p}_{i}(t)\|^{2}\phi\!\left(\mathbf{q}\right)\!\mathrm{d}\mathbf{q}$
$\!+\!\frac{1}{2}\!\sum_{i=1}^{n}\!\int_{Q}\frac{R_{i}\ddot{\lambda}(t)\lambda^{2}(t)-2R_{i}\dot{\lambda}^{2}(t)\lambda(t)}{\lambda^{4}(t)}h_{\lambda}(t,R_{i})$
$\quad\qquad\|\mathbf{q}\!-\!\mathbf{p}_{i}(t)\|^{2}\phi\!\left(\mathbf{q}\right)\!\mathrm{d}\mathbf{q}$
$\!-\!2\epsilon\!\sum_{i=1}^{n}\!\mathbf{W}^{\mathrm{T}}_{i}(t,\mathbf{P}(t))\frac{\partial\mathbf{W}_{i}(t,\mathbf{P}(t))}{\partial t}$
$-\!2\epsilon^{2}\!\sum_{i=1}^{n}\!\mathbf{W}^{\mathrm{T}}_{i}(t,\mathbf{P}(t))\sum_{j=1}^{N}\frac{\partial\mathbf{W}_{i}(t,\mathbf{P}(t))}{\partial\mathbf{p}_{j}(t)}\mathbf{W}_{j}(t,\mathbf{P}(t)).$ (36)
The first four terms in (IV-A) are of the form $\frac{R_{i}}{\lambda^{n}(t)}h_{\lambda}(R_{i})=\frac{R_{i}}{\lambda^{n}(t)}\exp(-R_{i}/\lambda(t))$, $n\in\{2,3,4\}$. Since $R_{i}$ takes integer values in $\{0,1,\dots,n-1\}$, if $R_{i}=0$, these four terms vanish identically. For $R_{i}\geq 1$, by definition $c(t)\triangleq 1/\lambda(t)$, we can deduce that $\lim_{t\to\infty}\frac{R_{i}}{\lambda^{n}(t)}\,h_{\lambda}(R_{i})=\lim_{C\to\infty}C^{n}e^{-C}=0$. Therefore, using Lemma 4, we can conclude that $\ddot{V}(\cdot)$ is uniformly bounded.
Since $V(\cdot)>0$ and $\dot{V}(\cdot)<0$, $V(\cdot)$ decreases monotonically to a finite positive limit. By using Barbalat’s lemma, $\dot{V}(\cdot)\to 0$, and from (33) and (23), we have $\lim_{t\to\infty}\mathbf{u}_{i}(t)=\lim_{t\to\infty}\mathbf{W}_{i}(t,\mathbf{P}(t))=0$, implying asymptotic convergence to the optimal coverage. $\blacksquare$